# Fixed-Point RNNs: Interpolating from Diagonal to Dense

### v2

@@ C.2 FP-Parametrization
In our design of FPMamba, we aim to minimize our interventions in the underlying architecture in order to showcase the adaptability of our proposed framework. Consequently, we do not modify the careful parameterization of $\boldsymbol{\lambda}$ and the weight-tied normalization factor $\Delta_{t}$ proposed in the original Mamba formulation, and instead rely on layer normalization to limit the Lipschitz constant of the Mamba function. Specifically, in the FPMamba model we redefine $\mathbf{B}_{t}$ and $\mathbf{C}_{t}$ as $\mathbf{B}_{t}^{\ell}=\mathbf{W}_{\mathbf{B}}^{y}y_{t-1}^{\ell-1}+\mathbf{W}_{\mathbf{B}}^{x}x_{t}$ and $\mathbf{C}_{t}=\mathbf{W}_{\mathbf{C}}^{y}y_{t-1}^{\ell-1}+\mathbf{W}_{\mathbf{C}}^{x}x_{t}$. The remaining components, namely the state transition matrix $\boldsymbol{\lambda}_{t}$ and the GLU component are parameterized identically to Mamba.
@@
The normalization is applied to the output of the model $y_{t}$ after each iteration. While in theory projecting the output onto the unit sphere does not guarantee a Lipschitz constant $<1$ , we observe that in practice, this helps with stabilizing the forward and backward pass of the fixed-point RNN framework. We attribute this observation to the fact that achieving a $>1$ Lipschitz constant requires the output of the RNN to become its additive inverse after an iteration, which rarely happens in practice.
@@ C.3 Parameterizing the mixers
[…]
@@
[…]
For the diagonal plus low rank and the Householder reflections mixers, we L2 normalize the vectors $\mathbf{u}_{it}$ to achieve the unit vector formulation. Note that this does not guarantee a contractive diagonal plus low rank structure, which is why the first variant of the channel mixers are excluded form our FPRNN experiments. For the Kronecker variant, we define the matrices $\mathbf{K}_{t}^{n}$ as symmetric and positive semi-definite using the Cholesky decompositon structure, and normalize them by their largest eigenvalues. The largest eigenvalue is found using the power iterations method, which we found to be much more efficient for small-scale matrices compared to the functions in the PyTorch framework provided for this purpose.
@@
Note that in order to scale the eigenvalue range of the mixers to negative eigenvalue, we can scale the coefficients $\alpha_{it}$ and $\mathbf{D}_{t}$ by a factor of $2$. This is similar to the formulation introduced in Grazzi et al. (2024) for the DeltaNet model.
C.4 Dependence on $h_{t-1}$ in theory
@@
We hypothesize that the dependence of the matrices $\boldsymbol{\lambda}_{t}$, $\mathbf{B}_{t}$, $\mathbf{C}_{t}$, and $\mathbf{Q}_{t}$ may provide a mechanism for the model to retain and manipulate positional information over the sequence. Jelassi et al. (2024) and Trockman et al. (2024) show that position embeddings could play a crucial role in copy tasks by acting similar to hashing keys in a hashing table. We extend their mechanistic approach to understand why two-layers of linear attention could need $h_{t-1}^{\ell-1}$ to generate appropriate position embeddings for the hashing mechanism.
@@
Specifically consider $y_{t}=\mathbf{C}_{t}^{\top}h_{t}$ with $h_{t}=h_{t-1}+\mathbf{B}_{t}x_{t}^{\top}$, assuming that a linear RNN with matrix-state can express linear attention by setting $\boldsymbol{\lambda}_{t}\approx\mathbf{1}~{}\forall t$. Upon receiving an input sequence $\{x_{1},x_{2},\dots,x_{\delta}\}$ of length $\delta$ followed by a delimiter element $x_{\mathbf{s}}$, the model is expected to copy the input sequence autoregressively, i.e. to start producing $\{x_{1},x_{2},\dots,x_{\delta}\}$ at output positions $\delta+1$ to $2\delta$. Following Arora et al. (2024), the second layer could use position embeddings as hashing keys to detect and copy each token. More concretely, if the first layer receives a sequence $\{x_{1},x_{2},\dots,x_{\delta},x_{\mathbf{s}},x_{1},x_{2},\dots,x_{\delta-1}\}$ of size $2\delta$ and augments it with shifted position embeddings $\{p_{i}\}_{i=1}^{\delta}$ to produce the hidden sequence $\{x_{1}+p_{1},x_{2}+p_{2},\dots,x_{\delta}+p_{\delta},x_{\mathbf{s}}+p_{1},x_{1}+p_{2},\dots,x_{\delta-1}+p_{\delta}\}$, then a second layer can act as a linear transformer and produce the sequence $\{x_{1},x_{2},\dots,x_{\delta}\}$ at output positions $\delta+1$ to $2\delta$. In the following, we focus on the conditions for the first layer to produce the shifted position embeddings.
@@
We start by assuming that the first layer has a skip-connection $y_{t}^{\top}=\mathbf{C}_{t}^{\top}h_{t}+x_{t}^{\top}$. In this case, the inputs can be augmented if the recurrence is able to produce shifted encodings $p_{t-\delta}=p_{t}$ for $\delta<t$ using $p_{t}^{\top}=\mathbf{C}_{t}^{\top}h_{t}$. This condition can be unrolled as
$\mathbf{C}_{t-\delta}^{\top}h_{t-\delta}\overset{!}{=}\mathbf{C}_{t}^{\top}h_{t-\delta}+\mathbf{C}_{t}^{\top}\sum_{\smash{\tau=t-\delta+1}}^{t}\mathbf{B}_{\tau}x_{\tau}^{\top}$ $\forall\delta<t.$
and is satisfied if the equations $\mathbf{C}_{t-\delta}^{\top}h_{t-\delta}\overset{!}{=}\mathbf{C}_{t}^{\top}h_{t-\delta}$ and
$\mathbf{C}_{t}^{\top}\sum_{\tau=t-\delta+1}^{t-1}\mathbf{B}_{\tau}x_{\tau}^{\top}\overset{!}{=}-\mathbf{C}_{t}^{\top}\mathbf{B}_{t}x_{t}^{\top}$
@@
hold. Such conditions could only be true if $\mathbf{B}_{t}$ and $\mathbf{C}_{t}$ are a function of the previous hidden state $h_{t-1}$ because they need to be able to retain information about $\{x_{i}\}_{i=t-\delta+1}^{t-1}$. While not an explicit mechanism for copying, this derivation provides insight into why a dependency on $h_{t-1}$ could be helpful.
@@ Appendix D Evaluation
D.1 Tasks
In this section, we provide experimental results for our proposed model. For a detailed summary of the experiment setup, please refer to App. D.2.
@@ D.1 Tasks
The task of tracking state in the alternating group on five elements ($A_{5}$) is one of the tasks introduced in (Merrill et al., 2024) to show that linear RNNs and SSMs cannot solve state-tracking problems. $A_{5}$ is the simplest subset of $S_{5}$, the word problem involving tracking the permutation of five elements. In these tasks, a model is presented with an initial state and a sequence of permutations. As the output, the model is expected to predict the state that results from applying the permutations to the initial state. Solving these task with an RNN requires either a dense transition matrix or the presence of non-linearity in the recurrence. It is therefore a good proxy to verify the state-tracking ability of FP-Mamba. In order to investigate the out-of-distribution generalization ability of the model, we train the model with a smaller train sequence length and evaluate for larger (more than ×$\times 3$) sequence lengths.
@@
We use the copy task (Jelassi et al., 2024) in order to assess the memory capabilities of FP-Mamba. In this task, the model is presented with a fixed-size sequence of elements, and expected to copy a subsequence of it after receiving a special token signaling the start of the copying process. In order to investigate the out-of-distribution generalization ability of the model, we train the models with sequence length $<50$, and assess the x2 length generalization following Jelassi et al. (2024) and Trockman et al. (2024).
The Chomsky Hierarchy
Following Grazzi et al. (2024), we also evaluate FP-Mamba on the remaining unsolved task of the Chomsky Hierarchy of language problems introduced by Beck et al. (2024). Specifically, we focus on the mod arithmetic task with brackets, for which the best performance reported so far according to Grazzi et al. (2024) is an accuracy of $0.2$. Following the setup of Grazzi et al. (2024), we train on sequence lengths $3$ to $40$ and report scaled accuracies on test sequences of lengths $40$ to $256$. For FP-Mamba, we use a $2$-layer model with $r=4$ reflections, i.e. the best performing model in the $A_{5}$ experiment.
D.2 Evaluation Details
In this section, we will provide our experiment setup for the state tracking, copying, and mod arithmetic tasks.
@@ D.2 Evaluation Details
We train all models for $5$ epochs, with a batch size of $512$, $3$ different random seeds, learning rate set to $0.0001$, weight decay set to $0.01$, gradient clipping $1.0$, and the AdamW optimizer (Loshchilov & Hutter, 2017). For the train data, we sample $16$M datapoints from all the possible permutations for a sequence length of $16$, and split the data with a ratio of $4$ to $1$ for train and validation samples. For the test data, we sample $500$k sequences of length $50$. We use the implementation and the hyperparameters provided by Merrill et al. (2024) both for data generation and train/test. We train the model for sequence length $16$ on the train sample, and evaluate for sequence lengths $2$ through $50$ on the test sample. Consequently, each epoch of training consists of $25428$ iterations, making the total number of iterations during training to be around 1.25M. Note that the likelihood of overlap between the train and test samples is negligible since exhaustive generation of samples in $S_{5}$ and $A_{5}$ at sequence length $k$ would amount to $60^{k}$ and $30^{k}$, respectively.
@@
We train all models for $10000$ iterations, batch size $128$, $3$ different random seeds, learning rate $0.00001$, weight decay $0.1$, gradient clipping $1.0$, the AdamW optimizer, and with linear learning rate decay after a $300$ iterations warmup. The data is sampled randomly at the start of the training/evaluation. We use a vocab size of $29$, a context length of $256$, and train the model for copy sequence length in the range $5$ to $50$, and evaluate for the range $5$ to $100$. we use the implementation and the hyperparameters provided by Jelassi et al. (2024).
@@
For the language modeling task, we use the implmentation provided by Ajroldi (2024). We use a batchsize of $16\times 4\times 4=256$, training on $4$ A100-80GB GPUs with $4$ accumulation steps, which is the batchsize used in the 2.5B setting in (Gu & Dao, 2023). The learning rate is optimized for the Mamba model ($0.004$) and train all models with this learning rate, with cosine warmup with $0.1$ steps. We use the AdamW optimizer with weight decay set to $0.1$ and $\beta_{1},\beta_{2}$ set to $0.9,0.95$.
@@ E.1 Language Modeling
Figure 8: The validation perplexity of the Mamba model vs. the FPMamba model for a Householder channel mixer of $r\in\{1,2\}$ reflections. Note that all of the hyperparameters of the models are identical for fair comparison.
In order to confirm the utility of the fixed-point framework in non-state-tracking settings, we performed an experiment on language modeling. Specifically, we compare the performance of a Mamba with an FPMamba, with the same hidden size ($768$) and number of layers ($12$). The settings are selected according to the 2.5B setup introduced in Gu & Dao (2023). We use a train subsample of the FineWeb dataset (Penedo et al., 2024) with 2B tokens, and a validation subsample with 200K tokens. We use a context length of $2048$ for our experiment. For the FPMamba model, we use the Householder mixer with $1$ and $2$ reflections. We report the validation perplexity in Fig. 8.
As we can observe, the fixed-point framework does introduce a significant improvement to the performance of the model on perplexity. However, we note that this improvement cannot be only attributed to the multi-layer hypothesis of implicit models (Giannou et al., 2023), as increasing the number of Householder reflections does seem to be improving the perplexity further. Furthermore, we point out the practicality of the setup, as we can observe in Fig. 7 that in the absence of a state-tracking problem, the number of fixed-point iterations seems to be independent of the sequence length, and instead hover in the $<10$ range. Finally, fixed-point iterations are not required in the backward bass and therefore only increase training time moderately.

### v3

@@ C.2 FP-Mamba Parametrization
In our design of FP-Mamba, we aim to minimize our interventions in the underlying architecture in order to showcase the adaptability of our proposed framework. Consequently, we do not modify the careful parameterization of $\boldsymbol{\lambda}$ and the weight-tied normalization factor $\Delta_{t}$ proposed in the original Mamba formulation, and instead rely on layer normalization to limit the Lipschitz constant of the Mamba function. Specifically, in the FP-Mamba model we redefine $\mathbf{b}_{t}$ and $\mathbf{c}_{t}$ as $\mathbf{b}_{t}^{\ell}=\mathbf{W}_{\mathbf{b}}^{\mathbf{y}}\mathbf{y}_{t-1}^{\ell-1}+\mathbf{W}_{\mathbf{b}}^{\mathbf{x}}\mathbf{x}_{t}$ and $\mathbf{c}_{t}=\mathbf{W}_{\mathbf{c}}^{\mathbf{y}}\mathbf{y}_{t-1}^{\ell-1}+\mathbf{W}_{\mathbf{c}}^{\mathbf{x}}\mathbf{x}_{t}$. The remaining components, namely the state transition matrix $\boldsymbol{\lambda}_{t}$ and the GLU component are parameterized identically to Mamba.
@@
The normalization is applied to the output of the model $\mathbf{y}_{t}$ after each iteration. While in theory projecting the output onto the unit sphere does not guarantee a Lipschitz constant <1<1 , we observe that in practice, this helps with stabilizing the forward and backward pass of the fixed-point RNN framework. We attribute this observation to the fact that achieving a >1>1 Lipschitz constant requires the output of the RNN to become its additive inverse after an iteration, which rarely happens in practice.
@@ C.3 Parameterizing the mixers
  • Diagonal Plus Low Rank: we define $\mathbf{u}_{it}^{\ell}=\text{SiLU}\left(\mathbf{W}_{\mathbf{u}_{i}}^{\mathbf{x}}\mathbf{x}_{t}+\mathbf{W}_{\mathbf{u}_{i}}^{\mathbf{y}}\mathbf{y}_{t-1}^{\ell-1}\right)$ and $\alpha_{it}=\sigma\left(\left(\mathbf{w}_{\alpha_{i}}^{\mathbf{x}}\right)^{\top}\mathbf{x}_{t}+(\mathbf{w}_{\alpha_{i}}^{\mathbf{y}})^{\top}\mathbf{y}_{t-1}^{\ell-1}+b_{\alpha_{i}}\right)$, where $\text{SiLU}(.)$ and $\sigma(.)$ are the SiLU and the sigmoid functions, respectively.
@@
  • Kronecker: we define $\mathbf{D}_{t}^{\ell,n}=\text{diag}\left(\sigma\left(\mathbf{W}_{\mathbf{D}^{n}}^{\mathbf{x}}\mathbf{x}_{t}+\mathbf{W}_{\mathbf{D}^{n}}^{\mathbf{y}}\mathbf{y}_{t-1}^{\ell-1}+b_{\mathbf{D}^{n}}\right)\right)$ and $\mathbf{K}_{t}^{n}=\text{mat}\left(\text{SiLU}\left(\mathbf{W}_{\mathbf{K}^{n}}^{\mathbf{x}}\mathbf{x}_{t}+\mathbf{W}_{\mathbf{K}^{n}}^{\mathbf{y}}\mathbf{y}_{t-1}^{\ell-1}+b_{\mathbf{K}^{n}}\right)\right)$ for $n=1,2$, where $\text{diag}(.)$ is the operator transforming a vector into a diagonal matrix, and $\text{mat}(.)$ is the operator transforming a size $d$ vector into a $\sqrt{d}\times\sqrt{d}$ matrix.
For the diagonal plus low rank and the Householder reflections mixers, we L2 normalize the vectors $\mathbf{u}_{it}$ to achieve the unit vector formulation. Note that this does not guarantee a contractive diagonal plus low rank structure, which is why the first variant of the channel mixers are excluded form our FP-RNN experiments. For the Kronecker variant, we define the matrices $\mathbf{K}_{t}^{n}$ as symmetric and positive semi-definite using the Cholesky decompositon structure, and normalize them by their largest eigenvalues. The largest eigenvalue is found using the power iterations method, which we found to be much more efficient for small-scale matrices compared to the functions in the PyTorch framework provided for this purpose.
@@
C.4 Dependence on $\mathbf{H}_{t-1}$ in theory
@@
We hypothesize that the dependence of the matrices $\boldsymbol{\lambda}_{t}$, $\mathbf{b}_{t}$, $\mathbf{c}_{t}$, and $\mathbf{Q}_{t}$ may provide a mechanism for the model to retain and manipulate positional information over the sequence. Jelassi et al. (2024) and Trockman et al. (2024) show that position embeddings could play a crucial role in copy tasks by acting similar to hashing keys in a hashing table. We extend their mechanistic approach to understand why two-layers of linear attention could need $\mathbf{H}_{t-1}^{\ell-1}$ to generate appropriate position embeddings for the hashing mechanism.
@@
Specifically consider $\mathbf{y}_{t}^{\top}=\mathbf{c}_{t}^{\top}\mathbf{H}_{t}$ with $\mathbf{H}_{t}=\mathbf{H}_{t-1}+\mathbf{b}_{t}\mathbf{x}_{t}^{\top}$, assuming that a linear RNN with matrix-state can express linear attention by setting $\boldsymbol{\lambda}_{t}\approx\mathbf{1}~\forall t$. Upon receiving an input sequence $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\delta}\}$ of length $\delta$ followed by a delimiter element $\mathbf{x}_{\mathbf{s}}$, the model is expected to copy the input sequence autoregressively, i.e. to start producing $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\delta}\}$ at output positions $\delta+1$ to $2\delta$. Following Arora et al. (2024), the second layer could use position embeddings as hashing keys to detect and copy each token. More concretely, if the first layer receives a sequence $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\delta},\mathbf{x}_{\mathbf{s}},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\delta-1}\}$ of size $2\delta$ and augments it with shifted position embeddings $\{\mathbf{p}_{i}\}_{i=1}^{\delta}$ to produce the hidden sequence $\{\mathbf{x}_{1}+\mathbf{p}_{1},\mathbf{x}_{2}+\mathbf{p}_{2},\dots,\mathbf{x}_{\delta}+\mathbf{p}_{\delta},\mathbf{x}_{\mathbf{s}}+\mathbf{p}_{1},\mathbf{x}_{1}+\mathbf{p}_{2},\dots,\mathbf{x}_{\delta-1}+\mathbf{p}_{\delta}\}$, then a second layer can act as a linear transformer and produce the sequence $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\delta}\}$ at output positions $\delta+1$ to $2\delta$. In the following, we focus on the conditions for the first layer to produce the shifted position embeddings.
@@
We start by assuming that the first layer has a skip-connection $\mathbf{y}_{t}^{\top}=\mathbf{c}_{t}^{\top}\mathbf{H}_{t}+\mathbf{x}_{t}^{\top}$. In this case, the model can augment the inputs with positional embeddings $\{\mathbf{p}_{i}\}_{i=1}^{\delta}$ if it is able to produce shifted encodings $\mathbf{p}_{t-\delta}=\mathbf{p}_{t}$ for $\delta<t$ using $\mathbf{p}_{t}^{\top}=\mathbf{c}_{t}^{\top}\mathbf{H}_{t}$. This condition can be unrolled as
$\mathbf{p}_{t-\delta}^{\top}=\mathbf{c}_{t-\delta}^{\top}\mathbf{H}_{t-\delta}\overset{!}{=}\mathbf{c}_{t}^{\top}\mathbf{H}_{t-\delta}+\mathbf{c}_{t}^{\top}\sum_{\smash{\tau=t-\delta+1}}^{t}\mathbf{b}_{\tau}\mathbf{x}_{\tau}^{\top}=\mathbf{p}_{t}^{\top}$ $\forall\delta<t.$
and is satisfied if the equations
$\mathbf{c}_{t-\delta}^{\top}\mathbf{H}_{t-\delta}\overset{!}{=}\mathbf{c}_{t}^{\top}\mathbf{H}_{t-\delta}$ and $\mathbf{c}_{t}^{\top}\sum_{\tau=t-\delta+1}^{t-1}\mathbf{b}_{\tau}\mathbf{x}_{\tau}^{\top}\overset{!}{=}-\mathbf{c}_{t}^{\top}\mathbf{b}_{t}\mathbf{x}_{t}^{\top}$
@@
hold. Such conditions could only be true if $\mathbf{b}_{t}$ and $\mathbf{c}_{t}$ are a function of the previous hidden state $\mathbf{H}_{t-1}$ because they need to be able to retain information about $\{\mathbf{x}_{i}\}_{i=t-\delta+1}^{t-1}$. While not an explicit mechanism for copying, this derivation provides insight into why a dependency on $\mathbf{H}_{t-1}$ could be helpful.
@@ Appendix D Evaluation
D.1 Task Descriptions
In this section, we provide task descriptions for the tasks used in the main text.
@@ D.1 Task Descriptions
The task of tracking state in the alternating group on five elements ($A_{5}$) is one of the tasks introduced in (Merrill et al., 2024) to show that linear RNNs and SSMs cannot solve state-tracking problems. $A_{5}$ is the simplest subset of $S_{5}$, the word problem involving tracking the permutation of five elements. In these tasks, a model is presented with an initial state and a sequence of permutations. As the output, the model is expected to predict the state that results from applying the permutations to the initial state. Solving these task with an RNN requires either a dense transition matrix or the presence of non-linearity in the recurrence. It is therefore a good proxy to verify the state-tracking ability of FP-Mamba. In order to investigate the out-of-distribution generalization ability of the model, we train the model with a smaller train sequence length and evaluate for larger (more than $\times 3$) sequence lengths.
@@
We use the copy task (Jelassi et al., 2024) in order to assess the memory capabilities of FP-Mamba. In this task, the model is presented with a fixed-size sequence of elements, and expected to copy a subsequence of it after receiving a special token signaling the start of the copying process. In order to investigate the out-of-distribution generalization ability of the model, we train the models with sequence length <50<50, and assess the $\times 2$ length generalization following Jelassi et al. (2024) and Trockman et al. (2024).
D.2 Experimental Details
In this section, we will provide our experiment setup for the state tracking, copying, and mod arithmetic tasks. The code is available at github.com/dr-faustus/fp-rnn.
@@ D.2 Experimental Details
We train all models for $5$ epochs, with a batch size of $512$, 33 different random seeds, learning rate set to $0.0001$, weight decay set to $0.01$, gradient clipping $1.0$, and the AdamW optimizer (Loshchilov & Hutter, 2017). For the train data, we sample 1616M datapoints from all the possible permutations for a sequence length of $16$, and split the data with a ratio of $4$ to $1$ for train and validation samples. For the test data, we sample 500500k sequences of length $50$. We use the implementation and the hyperparameters provided by Merrill et al. (2024) both for data generation and train/test. We train the model for sequence length $16$ on the train sample, and evaluate for sequence lengths $2$ through $50$ on the test sample. Consequently, each epoch of training consists of $25428$ iterations, making the total number of iterations during training to be around 1.25M. Note that the likelihood of overlap between the train and test samples is negligible since exhaustive generation of samples in $S_{5}$ and $A_{5}$ at sequence length $k$ would amount to $60^{k}$ and $30^{k}$, respectively.
@@
We train all models for $10000$ iterations, batch size $128$, 33 different random seeds, learning rate $0.00001$, weight decay $0.1$, gradient clipping $1.0$, the AdamW optimizer, and with linear learning rate decay after a $300$ iterations warmup. The data is sampled randomly at the start of the training/evaluation. We use a vocab size of $29$, a context length of $256$, and train the model for copy sequence length in the range $5$ to $50$, and evaluate for the range $5$ to $100$. we use the implementation and the hyperparameters provided by Jelassi et al. (2024).
@@
For the language modeling task, we use the implmentation provided by Ajroldi (2024). We use a batchsize of $16\times 4\times 4=256$, training on $4$ A100-80GB GPUs with $4$ accumulation steps, which is the batchsize used in the 2.5B setting in (Gu & Dao, 2024). The learning rate is optimized for the Mamba model ($0.004$) and train all models with this learning rate, with cosine warmup with $0.1$ steps. We use the AdamW optimizer with weight decay set to $0.1$ and $\beta_{1},\beta_{2}$ set to $0.9,0.95$.
@@
catbAbI
In this experiment, we use the setting provided by Schlag et al. (2021b). We optimize the learning rates on Mamba, and use the same learning rate to train FP-Mamba, which we found to be $5\times 10^{-4}$. We use a batch size of $256$, along with short convolutions, and $1$, 22, or $4$ layers. We set the maximum number of iterations $\ell_{\text{max}}$ to $100$.
D.3 Heuristics to reduce the number of fixed-point iterations
The second heuristic involves using a momentum-like update rule to accelerate the convergence of fixed-point iterations for certain sequences. Specifically, we observe that by setting the fixed-point update rule to $\mathbf{h}^{\ell+1}=\delta\cdot f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}^{\ell})+(1-\delta)\cdot\mathbf{h}^{\ell}$ for some $\delta\in[0,1]$, we can accelerate the convergence for certain sequences that are particularly slow to converge. Since this update rule can result in a biased approximation of the fixed-point, we implement a patience-based system that starts with $\delta=1$, and reduces the value of $\delta$ exponentially when the residues fail to improve.
@@ E.1 Language Modeling
Figure 8: The validation perplexity of the Mamba model vs. FP-Mamba-K and FP-Mamba-H with $r\in\{1,2\}$ reflections. Note that all of the hyperparameters of the models are identical for fair comparison.
In order to confirm the utility of the fixed-point framework in non-state-tracking settings, we performed an experiment on language modeling. Specifically, we compare the performance of a Mamba with an FP-Mamba, with the same hidden size ($768$) and number of layers ($12$). The settings are selected according to the 2.5B setup introduced in Gu & Dao (2024). We use a train subsample of the FineWeb dataset (Penedo et al., 2024) with 2B tokens, and a validation subsample with 200K tokens. We use a context length of $2048$ for our experiment. For the FP-Mamba model, we use the Householder mixer with $1$ and $2$ reflections. We report the validation perplexity in Fig. 8.
As we can observe, the fixed-point framework does introduce a significant improvement to the performance of the model on perplexity. However, we note that this improvement cannot be only attributed to the multi-layer hypothesis of implicit models (Giannou et al., 2023), as increasing the number of Householder reflections does seem to be improving the perplexity further. Furthermore, we point out the practicality of the setup, as we can observe in Fig. 7 that in the absence of a state-tracking problem, the number of fixed-point iterations seems to be independent of the sequence length, and instead hover in the <10<10 range. Finally, fixed-point iterations are not required in the backward bass and therefore only increase training time moderately.